OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

OpenStudy (anonymous):

y = 1/x has a horizontal asymptote of y = 0. Shift that up by 2/9: \[ \boxed {y = \dfrac {1}{x} + \dfrac {2}{9}}. \]

OpenStudy (anonymous):

so thatd be the answer.. the y=1/x+2/9

OpenStudy (anonymous):

??

OpenStudy (anonymous):

make the degree of the numerator equal to the degree of the denominator make the leading coefficient of the numerator 2 and the leading coefficient of the denominator 9

OpenStudy (anonymous):

wait.. what? im sorry i just got myself all confused

OpenStudy (anonymous):

@Ahaanomegas had a good answer if you add this up you get \[f(x)=\frac{2x+9}{9x}\] but anything similar would do for example \[f(x)=\frac{2x-10}{9x+17}\] or \[f(x)=\frac{2x^2+3x+1}{9x^2+5x-1}\] or anything where the degrees are the same and the leading coefficient of the numerator is 2 and the leading coefficient of the denominator is 9

OpenStudy (anonymous):

so either of those i could put for the answeR?

OpenStudy (anonymous):

??